OpenStudy (anonymous):

scientific notation 0.00045

OpenStudy (anonymous):

4.5*10^-4

OpenStudy (anonymous):

I guess u shud now write the answer for us to chekc whether u hv understood the method or not :))

OpenStudy (anonymous):

\[.00045=4.5\times .0001=4.5\times 10^{-4}\]

OpenStudy (anonymous):

why would it be a -4 ?

OpenStudy (anonymous):

because \[10^{-4}=\frac{1}{10^4}=\frac{1}{1000}=.0001\]

OpenStudy (anonymous):

because \[10^{-4}=\frac{1}{10^4}=\frac{1}{1000}=.0001\]

OpenStudy (anonymous):

because \[10^{-4}=\frac{1}{10^4}=\frac{1}{1000}=.0001\]

OpenStudy (anonymous):

because \[10^{-4}=\frac{1}{10^4}=\frac{1}{1000}=.0001\]

OpenStudy (anonymous):

because \[10^{-4}=\frac{1}{10^4}=\frac{1}{1000}=.0001\]

OpenStudy (anonymous):

sorry computer is messing up